# Causally Reliable Concept Bottleneck Models

### v3

@@ 1 Introduction
In recent years, interpretable neural models have become more popular, achieving performance similar to powerful opaque Deep Neural Networks (DNNs) (alvarez2018towards; chen2019looks; chen2020concept). Among these, Concept Bottleneck Models (CBMs) (koh2020; zarlenga2022concept; yuksekgonul2022post; barbiero2023interpretable) guarantee high expressivity and interpretability by enforcing DNNs to reason through a layer of high-level, human-interpretable variables called concepts (e.g., the “color” and “shape” of an object) (kim2018interpretability; achtibat2023attribution; fel2023craft). In CBMs, a neural encoder first maps the raw input to concepts, forming a semantically transparent intermediate representation that is used by a simple decoder for downstream predictions. Beyond transparency, this design allows human experts to intervene on mispredicted concepts at test time to improve downstream task predictions (espinosa2024learning).
However, like standard DNN architectures, CBMs remain pure associative models (Pearl2019): their decision-making process reflects statistical correlations within the data rather than real-world causal mechanisms. As a result, they fail to distinguish between spurious correlations and true causal relationships. Recognizing this distinction is fundamental to achieving a reliable scientific understanding, supporting causal reasoning for intervention (Pearl2009; peters2017elements), enabling robust generalization under distributional shifts, and the implementation of fairness constraints (scholkopf2021toward; wang2022out).
@@ 2 Preliminaries
CBMs (koh2020) are interpretable-by-design architectures that explain their predictions using high-level interpretable variables called concepts. Standard CBMs decompose prediction into two stages: a neural encoder maps the input $X$ to a set of intermediate concepts $\mathcal{V}=\{V_{i}\}_{i=1}^{C}$, and a decoder predicts the target $Y$ from $\mathcal{V}$. This yields:
@@
Concept Embedding Models (CEMs) (zarlenga2022concept) enhance CBMs by pairing concepts with high-dimensional embeddings of the form $P(\mathcal{U}\mid\mathcal{V},X)$, where $\mathcal{U}=\{U_{i}\}_{i=1}^{C}$. These embeddings are provided to the decoder to predict the target variable $Y$, enabling the model to achieve performance comparable to standard DNN approaches while maintaining semantic interpretability. Critically, traditional decoders rely on a bipartite structure assumption, wherein all concepts are treated as direct causes of the target, e.g., $Y=f(V_{1},\dots,V_{C})$ for CBMs. This assumption is often overly simplistic for real-world problems. Bringing the reasoning of concept-based architectures closer to real-world mechanisms constitutes the main focus of this work.
@@
A model $\mathcal{M}$ is causally reliable w.r.t. a target phenomenon $T$ if and only if the structure of $\mathcal{M}$’s decision-making process is consistent with the causal mechanisms underlying $T$ (Termine2023Causal). Although state-of-the-art DNNs and concept-based models offer high expressivity, they lack causal reliability.
@@
The standard framework for modeling causal mechanisms is the structural causal model (SCM) (bareinboim2022pearl). An SCM $\mathcal{M}$ is a tuple $\langle\mathcal{V},\,\mathcal{U},\,\mathcal{F},\,P\rangle$, where:
@@
Each SCM can be associated with a graphical representation in which nodes correspond to the variables, and edges encode the functional relationships specified by $\mathcal{F}$. Here, we focus on SCMs whose associated graph is a directed acyclic graph (DAG) (pearl1995causal; zaffalon2020structural). In most cases, the underlying DAG is unknown and must be inferred from observational data, a process known as causal discovery (peters2017elements; zanga2022). However, methods based solely on observational data cannot generally guarantee the identification of a unique DAG (peters2017elements). The set of candidate DAGs can be refined by incorporating additional information, which we refer to as background knowledge (andrews2020completeness; abdulaal2023causal). This can be drawn from a range of sources, such as human experts, structured repositories of information (e.g., domain ontologies), or “unstructured” samples of information (e.g., scientific papers or other documentation).
@@ 3 Related works
Traditional concept-based architectures impose a strict bipartite structure in which concept neuron activations are assumed to directly cause task outputs (koh2020; yuksekgonul2022post; kim2023probabilistic; oikarinen2023; yang2023language; barbiero2023interpretable; vandenhirtz2024stochastic). This strong, often unrealistic assumption can lead to misleading explanations. For example, attributing a lung cancer diagnosis to both a ‘cough’ and ‘smoker’ concept could risk the false interpretation that reducing coughing could reduce cancer risk. Moreover, most CBMs assume independence among concepts, which is unrealistic, as it ignores natural co-occurrences (e.g., ‘smoke’ and ‘fire’) and prevents improvements in one concept from propagating to related concepts during interventions. Stochastic CBM (SCBM) (vandenhirtz2024stochastic) and Concept Graph Models (dominici2025causal) attempt to relax this assumption. However, these approaches capture only associations rather than causal relations, making it vulnerable to spurious correlations in the data. To date, no methodology exists for structuring the concept bottleneck according to a reliable causal model.
Recent approaches like DiConStruct (moreira2024diconstruct), aim to improve this aspect by generating causal graphs linking concepts to opaque DNN predictions. However, DiConStruct is a post-hoc method that may misalign with the original DNN’s outputs and relies solely on observational data, neglecting background knowledge and resulting in under-determined causal structures. Other architectures, such as Neural Causal Models (ke2019learning) and Neural Causal Abstractions (xia2024neural), impose even stronger assumptions, requiring access to either the true causal graph or a low-resolution structural causal model, which are impractical in many cases.
@@ 4 Method
Figure 2: Overview of the C2BM fully automated pipeline. The pipeline consists of three key blocks: (i) discovery and labeling of the relevant variables $\mathcal{V}$ from background knowledge; (ii) discovery of the causal graph by integrating data and background knowledge; (iii) a C2BM model, comprising a neural encoder and an adaptively parametrized SCM. Once the model is trained, it can support queries about any endogenous variable (e.g., predicting dyspea).
@@ 4.2 Model instantiation
Concepts and causal graph constitute an input for C2BM, which could remain agnostic to how they are obtained, e.g., provided by human experts. Notably, alternative or novel approaches may be employed, provided they solve the same problems (loula2025syntactic).
@@ 4.2.1 Concept discovery and labeling
In the CBM community, automated concept discovery and labeling using Large Language Models (LLMs) has become a standard solution when human supervision is unavailable (oikarinen2023; yang2023language; srivastava2024vlg; yamaguchi2025explanation). In our implementation, we follow the label-free CBM approach from oikarinen2023, where concepts are discovered by querying an LLM for those most relevant to the task. We then apply a filtering procedure to retain only concepts that meet criteria such as brevity, distinctiveness (i.e., not too similar to each other or the target), and presence in the training data.
Once $\mathcal{V}$ are selected, we label the dataset with variable annotations $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{v}_{i})\}_{i=1}^{N}$ to supervise concept learning. To do so, we adopt a strategy similar to oikarinen2023, leveraging a pre-trained contrastive vision-language model, such as CLIP (radford2021learning). This projects both data samples and the discovered concept names into a shared embedding space and computes their alignment to generate concept labels. Full implementation details are provided in App. B.
@@ 4.2.2 Causal graph discovery
Let $\mathbf{G}^{*}$ be the true, unknown, graph over a set of variables $\mathcal{V}$ from which a dataset $\mathcal{D}$ was generated. The causal discovery problem consists in recovering $\mathbf{G}^{*}$ from the observed dataset $\mathcal{D}$ (zanga2022).
As anticipated in Sec. 2, a promising direction for addressing this problem is to combine standard causal discovery algorithms with knowledge-base querying. In our pipeline, we focus on a well-known class of methods that recover an equivalence class of graphs from data, referred to as the Markov equivalence class (MEC) (Spirtes2001; Pearl2009; zanga2022).This class can be compactly represented as a Completed Partially Directed Acyclic Graph (CPDAG), an extension of a DAG where edges remain unoriented when there is insufficient evidence in the data to infer causal direction. This is a desirable property as it prevents incorrect (spurious) orientations based on the data alone. Specifically, we apply the Greedy Equivalence Search (GES) (chickering2002) algorithm, which we found performed well empirically (see App. G.3). Then, we leverage a pre-trained LLM to assess each undirected edge in the CPDAG, orienting the ones corresponding to true causal relationships while discarding those originating from spurious correlations. To improve the robustness and generalizability of this approach, we pair the LLM with a Retrieval Augmented Generation (RAG) technique, which is known to reduce hallucinations and can provide problem-specific knowledge. To further improve robustness and performance, we repeat each query 10 times, selecting the most frequent outcome (wang2022self). Implementation details are provided in App. C.
@@ 4.3 Structural equations and model training
where $\textsf{PA}_{i}$} denotes the set of endogenous parents of $V_{i}$ [^2] We refer to $V_{i}$ as a variable to allow for a more general formalization. In a classification setting, concepts assume categorical values; hence $V_{i}$ represents the probability of a concept state activation.. For the parameters $\bm{\theta}_{f_{i}}$}, we do not learn a single parameterization; instead, these are adaptively inferred for each different realization of $X$, by a hypernetwork $\mathbf{r}(\cdot)$ (ha2017hypernetworks; barbiero2023interpretable; debot2024interpretable), based on the values of the exogenous variables $\mathcal{U}$ and the graph connectivity. In our implementation, we consider separate hypernetworks $\mathbf{r}_{i}(\cdot)$ (e.g., independent DNNs), each taking as input a separate exogenous variable:
@@
The design in Eq. 3 and 4 improves both interpretability and expressivity. To aid (mechanistic) interpretability, structural equations take a linear form (Eq. 3): the value of each is a weighted linear combination of its parents. The adaptive re-parameterization of the equation’s weights performed by the hypernetwork $\mathbf{r}(\cdot)$ allows the model to also approximate non-linear relationships among endogenous variables (see App. D, which also includes a proof that C2BM is a universal approximator, regardless of the underlying causal graph). This idea is in line with existing literature on interpretability (ribeiro2016should; alvarez2018towards) and concept-based methods (barbiero2023interpretable)  [^3] This idea also aligns with balke1994counterfactual and zaffalon2020, where exogenous variables are used to represent relationships between endogenous variables when the structural equations are unknown.
@@
We clarify that we do not claim to identify the true structural functions. Instead, C2BM numerically approximates the outcomes as if they were generated by the underlying (unknown) structural equations. This approximation, along with C2BM’s DAG, is sufficient to compute reliable interventions, which is a core objective in the concept-based community (Poeta2023; steinmann2024learning).
@@ 5 Experimental evaluations
The considered datasets include both synthetic and real-world benchmarks. As synthetic datasets, we sample $10^{4}$ points from each of the five following discrete Bayesian networks available from the bnlearn repository (scutari2010): Asia (asia), Sachs (sachs2005causal), Insurance (binder1997), Alarm (alarm), and Hailfinder (abramson1996). We include cMNIST, a variant of the original dataset (lecun2010) in which the image data are colored according to custom rules. Additionally, we consider three real-world datasets: CelebA (liu2015), a facial recognition dataset labeled with different binary facial attributes; CUBC, a custom version of the original bird image dataset (he2019fine) from which we select a subset of concepts and define new ones to introduce deeper causal relationships; Siim-Pneumothorax (you2023), containing chest X-ray images annotated with a single label indicating the presence of pneumothorax, without additional concepts or their annotations. To generate them, we follow the label-free approach outlined in Sec. 4.2.1. Exhaustive details on all datasets are given in App. E.
The performance of the proposed pipeline is investigated alongside an opaque neural baseline predicting the task variable only (OpaqNN) and established state-of-the-art (SOTA) concept-based architectures, namely: CBM (koh2020), with linear and non-linear decoder; CEM (zarlenga2022concept); and SCBM (vandenhirtz2024stochastic). Hyperparameters have been selected via an independent search for each dataset–model pair based on performance on the validation set. Further details on each model’s architecture and hyperparameters can be found in App. F. Note that all baselines, except for OpaqNN, provide concept-based explanations for their predictions and allow concept interventions at test-time. This excludes architectures such as Self-Explainable Neural Networks (alvarez2018towards) and Concept Whitening (chen2020concept) as they do not offer a clear mechanism for intervening on their concept bottlenecks. We also excluded other CBM baselines such as Probabilistic CBMs (kim2023probabilistic), Post-hoc CBMs (yuksekgonul2022post), Label-free CBMs (oikarinen2023; yang2023language), as all of them share the same limitation of vanilla CBMs and CEMs: the causal graph is fixed and bipartite. Python code for reproducing all experiments is provided alongside the submission as supplementary material.

### v4

@@ 1 Introduction
In recent years, interpretable neural models have become more popular, achieving performance similar to powerful opaque Deep Neural Networks (DNNs) (Alvarez Melis and Jaakkola, 2018; Chen et al., 2019, 2020). Among these, Concept Bottleneck Models (CBMs) (Koh et al., 2020; Zarlenga et al., 2022; Yuksekgonul et al., 2022; Barbiero et al., 2023) guarantee high expressivity and interpretability by enforcing DNNs to reason through a layer of high-level, human-interpretable variables called concepts (e.g., the “color” and “shape” of an object) (Kim et al., 2018; Achtibat et al., 2023; Fel et al., 2023). In CBMs, a neural encoder first maps the raw input to concepts, forming a semantically transparent intermediate representation that is used by a simple decoder for downstream predictions. Beyond transparency, this design allows human experts to intervene on mispredicted concepts at test time to improve downstream task predictions (Espinosa Zarlenga et al., 2024).
However, like standard DNN architectures, CBMs remain pure associative models (Pearl, 2019): their decision-making process reflects statistical correlations within the data rather than real-world causal mechanisms. As a result, they fail to distinguish between spurious correlations and true causal relationships. Recognizing this distinction is fundamental to achieving a reliable scientific understanding, supporting causal reasoning for intervention (Pearl, 2009; Peters et al., 2017), enabling robust generalization under distributional shifts, and the implementation of fairness constraints (Schölkopf et al., 2021; Wang et al., 2022).
@@ 2 Preliminaries
CBMs (Koh et al., 2020) are interpretable-by-design architectures that explain their predictions using high-level interpretable variables called concepts. Standard CBMs decompose prediction into two stages: a neural encoder maps the input $X$ to a set of intermediate concepts $\mathcal{V}=\{V_{i}\}_{i=1}^{C}$, and a decoder predicts the target $Y$ from $\mathcal{V}$. This yields:
@@
Concept Embedding Models (CEMs) (Zarlenga et al., 2022) enhance CBMs by pairing concepts with high-dimensional embeddings of the form $P(\mathcal{U}\mid\mathcal{V},X)$, where $\mathcal{U}=\{U_{i}\}_{i=1}^{C}$. These embeddings are provided to the decoder to predict the target variable $Y$, enabling the model to achieve performance comparable to standard DNN approaches while maintaining semantic interpretability. Critically, traditional decoders rely on a bipartite structure assumption, wherein all concepts are treated as direct causes of the target, e.g., $Y=f(V_{1},\dots,V_{C})$ for CBMs. This assumption is often overly simplistic for real-world problems. Bringing the reasoning of concept-based architectures closer to real-world mechanisms constitutes the main focus of this work.
@@
A model $\mathcal{M}$ is causally reliable w.r.t. a target phenomenon $T$ if and only if the structure of $\mathcal{M}$’s decision-making process is consistent with the causal mechanisms underlying $T$ (Termine and Primiero, 2024). Although state-of-the-art DNNs and concept-based models offer high expressivity, they lack causal reliability.
@@
The standard framework for modeling causal mechanisms is the structural causal model (SCM) (Bareinboim et al., 2022). An SCM $\mathcal{M}$ is a tuple $\langle\mathcal{V},\,\mathcal{U},\,\mathcal{F},\,P\rangle$, where:
@@
Each SCM can be associated with a graphical representation in which nodes correspond to the variables, and edges encode the functional relationships specified by $\mathcal{F}$. Here, we focus on SCMs whose associated graph is a directed acyclic graph (DAG) (Pearl, 1995; Zaffalon et al., 2020b). In most cases, the underlying DAG is unknown and must be inferred from observational data, a process known as causal discovery (Peters et al., 2017; Zanga et al., 2022). However, methods based solely on observational data cannot generally guarantee the identification of a unique DAG (Peters et al., 2017). The set of candidate DAGs can be refined by incorporating additional information, which we refer to as background knowledge (Andrews et al., 2020; Abdulaal et al., 2023). This can be drawn from a range of sources, such as human experts, structured repositories of information (e.g., domain ontologies), or “unstructured” samples of information (e.g., scientific papers or other documentation).
@@ 3 Related works
Traditional concept-based architectures impose a strict bipartite structure in which concept neuron activations are assumed to directly cause task outputs (Koh et al., 2020; Yuksekgonul et al., 2022; Kim et al., 2023; Oikarinen et al., 2023; Yang et al., 2023; Barbiero et al., 2023; Vandenhirtz et al., 2024). This strong, often unrealistic assumption can lead to misleading explanations. For example, attributing a lung cancer diagnosis to both a ‘cough’ and ‘smoker’ concept could risk the false interpretation that reducing coughing could reduce cancer risk. Moreover, most CBMs assume independence among concepts, which is unrealistic, as it ignores natural co-occurrences (e.g., ‘smoke’ and ‘fire’) and prevents improvements in one concept from propagating to related concepts during interventions. Stochastic CBM (SCBM) (Vandenhirtz et al., 2024) and Concept Graph Models (Dominici et al., 2025) attempt to relax this assumption. However, these approaches capture only associations rather than causal relations, making it vulnerable to spurious correlations in the data. To date, no methodology exists for structuring the concept bottleneck according to a reliable causal model.
Recent approaches like DiConStruct (Moreira et al., 2024), aim to improve this aspect by generating causal graphs linking concepts to opaque DNN predictions. However, DiConStruct is a post-hoc method that may misalign with the original DNN’s outputs and relies solely on observational data, neglecting background knowledge and resulting in under-determined causal structures. Other architectures, such as Neural Causal Models (Ke et al., 2019) and Neural Causal Abstractions (Xia and Bareinboim, 2024), impose even stronger assumptions, requiring access to either the true causal graph or a low-resolution structural causal model, which are impractical in many cases.
@@ 4 Method
Figure 2: Overview of the C2BM fully automated pipeline. The pipeline consists of three key blocks: (i) Concept discovery: discovery and labeling of the relevant variables $\mathcal{V}$ from background knowledge; (ii) Causal graph discovery: discovery of the causal graph by integrating data and background knowledge; (iii) the C2BM model, comprising a neural encoder and an adaptively parametrized SCM. Once the model is trained, it can support forward and interventional queries about any endogenous variable (e.g., predicting dyspea).
@@ 4.2 Model instantiation
Concepts and causal graph constitute an input for C2BM, which could remain agnostic to how they are obtained, e.g., provided by human experts. Notably, alternative or novel approaches may be employed, provided they solve the same problems (Loula et al., 2025).
@@ 4.2.1 Concept discovery and labeling
In the CBM community, automated concept discovery and labeling using Large Language Models (LLMs) has become a standard solution when human supervision is unavailable (Oikarinen et al., 2023; Yang et al., 2023; Srivastava et al., 2024; Yamaguchi and Nishida, 2025). In our implementation, we follow the label-free CBM approach from Oikarinen et al. (2023), where concepts are discovered by querying an LLM for those most relevant to the task. We then apply a filtering procedure to retain only concepts that meet criteria such as brevity, distinctiveness (i.e., not too similar to each other or the target), and presence in the training data.
Once $\mathcal{V}$ are selected, we label the dataset with variable annotations $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{v}_{i})\}_{i=1}^{N}$ to supervise concept learning. To do so, we adopt a strategy similar to Oikarinen et al. (2023), leveraging a pre-trained contrastive vision-language model, such as CLIP (Radford et al., 2021). This projects both data samples and the discovered concept names into a shared embedding space and computes their alignment to generate concept labels. Full implementation details are provided in App. B.
@@ 4.2.2 Causal graph discovery
Let $\mathbf{G}^{*}$ be the true, unknown, graph over a set of variables $\mathcal{V}$ from which a dataset $\mathcal{D}$ was generated. The causal discovery problem consists in recovering $\mathbf{G}^{*}$ from the observed dataset $\mathcal{D}$ (Zanga et al., 2022).
As anticipated in Sec. 2, a promising direction for addressing this problem is to combine standard causal discovery algorithms with knowledge-base querying. In our pipeline, we focus on a well-known class of methods that recover an equivalence class of graphs from data, referred to as the Markov equivalence class (MEC) (Spirtes et al., 2001; Pearl, 2009; Zanga et al., 2022).This class can be compactly represented as a Completed Partially Directed Acyclic Graph (CPDAG), an extension of a DAG where edges remain unoriented when there is insufficient evidence in the data to infer causal direction. This is a desirable property as it prevents incorrect (spurious) orientations based on the data alone. Specifically, we apply the Greedy Equivalence Search (GES) (Chickering, 2002) algorithm, which we found performed well empirically (see App. G.3). Then, we leverage a pre-trained LLM to assess each undirected edge in the CPDAG, orienting the ones corresponding to true causal relationships while discarding those originating from spurious correlations. To improve the robustness and generalizability of this approach, we pair the LLM with a Retrieval Augmented Generation (RAG) technique, which is known to reduce hallucinations and can provide problem-specific knowledge. To further improve robustness and performance, we repeat each query 10 times, selecting the most frequent outcome (Wang et al., 2023). Implementation details are provided in App. C.
@@ 4.3 Structural equations and model training
where $\textsf{PA}_{i}$} denotes the set of endogenous parents of $V_{i}$ [^2] We refer to $V_{i}$ as a variable to allow for a more general formalization. In a classification setting, concepts assume categorical values; hence $V_{i}$ represents the probability of a concept state activation.. For the parameters $\bm{\theta}_{f_{i}}$}, we do not learn a single parameterization; instead, these are adaptively inferred for each different realization of $X$, by a hypernetwork $\mathbf{r}(\cdot)$ (Ha et al., 2017; Barbiero et al., 2023; Debot et al., 2024), based on the values of the exogenous variables $\mathcal{U}$ and the graph connectivity. In our implementation, we consider separate hypernetworks $\mathbf{r}_{i}(\cdot)$ (e.g., independent DNNs), each taking as input a separate exogenous variable:
@@
The design in Eq. 3 and 4 improves both interpretability and expressivity. To aid (mechanistic) interpretability, structural equations take a linear form (Eq. 3): the value of each is a weighted linear combination of its parents. The adaptive re-parameterization of the equation’s weights performed by the hypernetwork $\mathbf{r}(\cdot)$ allows the model to also approximate non-linear relationships among endogenous variables (see App. D, which also includes a proof that C2BM is a universal approximator, regardless of the underlying causal graph). This idea is in line with existing literature on interpretability (Ribeiro et al., 2016; Alvarez Melis and Jaakkola, 2018) and concept-based methods (Barbiero et al., 2023)  [^3] This idea also aligns with Balke and Pearl (1994) and Zaffalon et al. (2020a), where exogenous variables are used to represent relationships between endogenous variables when the structural equations are unknown.
@@
We clarify that we do not claim to identify the true structural functions. Instead, C2BM numerically approximates the outcomes as if they were generated by the underlying (unknown) structural equations. This approximation, along with C2BM’s DAG, is sufficient to compute reliable interventions, which is a core objective in the concept-based community (Poeta et al., 2023; Steinmann et al., 2024).
@@ 5 Experimental evaluations
The considered datasets include both synthetic and real-world benchmarks. As synthetic datasets, we sample $10^{4}$ points from each of the five following discrete Bayesian networks available from the bnlearn repository (Scutari, 2010): Asia (Lauritzen and Spiegelhalter, 1988), Sachs (Sachs et al., 2005), Insurance (Binder et al., 1997), Alarm (Beinlich et al., 1989), and Hailfinder (Abramson et al., 1996). We include cMNIST, a variant of the original dataset (LeCun et al., 2010) in which the image data are colored according to custom rules. Additionally, we consider three real-world datasets: CelebA (Liu et al., 2015), a facial recognition dataset labeled with different binary facial attributes; CUBC, a custom version of the original bird image dataset (He and Peng, 2019) from which we select a subset of concepts and define new ones to introduce deeper causal relationships; Siim-Pneumothorax (You et al., 2023), containing chest X-ray images annotated with a single label indicating the presence of pneumothorax, without additional concepts or their annotations. To generate them, we follow the label-free approach outlined in Sec. 4.2.1. Exhaustive details on all datasets are given in App. E.
The performance of the proposed pipeline is investigated alongside an opaque neural baseline predicting the task variable only (OpaqNN) and established state-of-the-art (SOTA) concept-based architectures, namely: CBM (Koh et al., 2020), with linear and non-linear decoder; CEM (Zarlenga et al., 2022); and SCBM (Vandenhirtz et al., 2024). Hyperparameters have been selected via an independent search for each dataset–model pair based on performance on the validation set. Further details on each model’s architecture and hyperparameters can be found in App. F. Note that all baselines, except for OpaqNN, provide concept-based explanations for their predictions and allow concept interventions at test-time. This excludes architectures such as Self-Explainable Neural Networks (Alvarez Melis and Jaakkola, 2018) and Concept Whitening (Chen et al., 2020) as they do not offer a clear mechanism for intervening on their concept bottlenecks. We also excluded other CBM baselines such as Probabilistic CBMs (Kim et al., 2023), Post-hoc CBMs (Yuksekgonul et al., 2022), Label-free CBMs (Oikarinen et al., 2023; Yang et al., 2023), as all of them share the same limitation of vanilla CBMs and CEMs: the causal graph is fixed and bipartite. Python code for reproducing all experiments is provided alongside the submission as supplementary material.